Simplify the following Boolean function: F(A,B,C) = BC + AC + ABC with dont care terms = ABC + ABC: O A+C AB+C O AC O
Answer #1

F(A, B, C) = BITÀCABC In Canonical form FCA,B,C) - (A+ĀJ B+ ĀL BPBDC +ABCg = ABT + Ā B C + ĀBC+ĀDc+ABC =&mco, 1,3,4,5) DontIF (A, B, C) = Btc. I Hence the mainimize expression - FIA,B,C) = Btc or A + B IF (A, B, C) = None of the options are . Corre